# Quantum Strategies to Overcome Classical Multiplexing Limits

### v1

@@ I Introduction
Throughout this paper, our primary figure of merit will be the improvement introduced by multiplexing, which we define to be
@@
for a $s$-qubit protocol, with $R_{s|F}(M)$ the maximum rate achievable at a desired minimum final state fidelity $F$, $M$ the abstract multiplexing resource, and all functions implicitly depending on many experimental parameters. For single qubit protocols we drop the subscript and denote the advantage $m$. Physically, this quantity $m_{s}$ defined in Eq. 2 is precisely how many copies of the un-multiplexed protocol would need to run in parallel to give the same rate as one implementation of the multiplexed protocol, subject to standard statistical assumptions (i.i.d time-independent statistics).
@@ II.1 Example: Temporal Multiplexing for Remote State Preparation with Weak Coherent Pulses
Assuming a sufficiently stringent fidelity is chosen such that $w\gg s$, for $p\rightarrow 0$ we arrive at the tighter bound on the improvement factor $m_{s}\leq M^{s}$. In this limit, we would need $M^{s}$ independent (that is, not sharing qubits or resources) copies of an un-multiplexed experimental implementation of this RSP protocol to yield the same rate of $s$-qubit delivery that we can achieve with a single multiplexed implementation making use of $M$ temporal modes. This superlinearity illustrates the resource efficiency of semiclassical temporal multiplexing for many-qubit protocols.
@@ IV.2 Toy Model for $s=2$: Analytic Solutions
Taking the low-success limit such that $P_{sc}\ll 1/n_{e/o}$, but taking $P_{ss}$ large such that $1-(1-P_{ss})^{n_{e}}$ is close to unity, we can divide Eqs. 33-34 by Eq. 33 with $M=1$ (the unmultiplexed case) to yield the $s=2$ multiplexing improvement factors for our two strategies
@@ VI Acknowledgements
We gratefully acknowledge Mariagrazia Iuliano, Bart van der Vecht, Thomas Beachamp, Álvaro Iñesta, Harold Ollivier, Maxime Garnier, Tracy Northrup, Ronald Hanson, and Benedikt Tissot for stimulating and illuminating conversations throughout the course of this project, as well as the membership of the Quantum Internet Alliance for their broad support.
@@ Appendix A Derivation of Semiclassical Limit to Multiplexing
Here we derive rigorously the semiclassical upper-bounds to the multiplexing advantage introduced in section II. We begin with the derivation of our first limit Eq. 6: the multiplexing advantage to the single-qubit rate to prepare or transmit a qubit of fidelity $F$, defined
$m=\frac{R_{p_{M}|F}(M)}{R_{p|F}(1)}$ (41)
where to be explicit we have introduced a probabilities $p_{M}/p$ for each individual attempt to succeed with/without multiplexing. Since $p_{M}\leq p$ (otherwise, we would not do it), we will take equality $p_{M}=p$ to give an upper limit to $m$. We assume that the classical multiplexing does not negatively impact the maximum achievable fidelity for the protocol, as is the case e.g. for temporal multiplexing.
We now recast our un-multiplexed rate $R_{p|F}(1)$ in terms of expectation values $\tau_{p}$ the time to achieve a first success, which is a random variable. In this way, $R_{p|F}(1)=\frac{1}{\mathbb{E}(\tau_{p})(\tau_{1}+\tau_{c})}$, with $\tau_{1}$ the duration of a single attempt’s propogation time and $\tau_{c}$ the duration of any additional classical communication needed to complete the attempt.

### v2

@@ I Introduction
Throughout this paper, we will study the improvement introduced by multiplexing to the rate of execution of a quantum protocol. We quantify this by the multiplexing improvement factor $m_{s}$: the ratio of the multiplexed $s$-qubit protocol execution rate to the un-multiplexed $s$-qubit protocol execution rate. Mathematically, we define
@@
for a $s$-qubit protocol, with $R_{s|F}(M)$ the maximum rate achievable at a desired minimum final state fidelity $F$, $M$ the abstract multiplexing resource, and all functions implicitly depending on many experimental parameters. For single qubit protocols we drop the subscript and denote the advantage $m$.
@@ II.1 Example: Temporal Multiplexing for Remote State Preparation with Weak Coherent Pulses
Assuming a sufficiently stringent fidelity is chosen such that $w\gg s$, for $p\rightarrow 0$ we arrive at the tighter bound on the improvement factor $m_{s}\leq M^{s}$. In this limit, the overall process is Poissonian and we can give a nice physical interpretation; we would need $M^{s}$ independent (that is, not sharing qubits or resources) copies of an un-multiplexed experimental implementation of this RSP protocol to yield the same rate of $s$-qubit delivery that we can achieve with a single multiplexed implementation making use of $M$ temporal modes. This superlinearity illustrates the resource efficiency of semiclassical temporal multiplexing for many-qubit protocols.
@@ IV.2 Toy Model for $s=2$: Analytic Solutions
Taking the low-success limit such that $P_{sc}\ll 1/n_{e/o}$, but taking $P_{ss}$ large such that $1-(1-P_{ss})^{n_{e}}$ is close to unity, we can divide Eqs. 33-34 by Eq. 33 with $M=1$ (the un-multiplexed case) to yield the $s=2$ multiplexing improvement factors for our two strategies
@@ VI Acknowledgements
We gratefully acknowledge assistance provided by Bethany Davies and Thomas Beauchamp with our analysis. We are also grateful to Mariagrazia Iuliano, Bart van der Vecht, Álvaro Iñesta, Harold Ollivier, Maxime Garnier, Tracy Northrup, Ronald Hanson, and Benedikt Tissot for stimulating and illuminating discussion throughout the course of this project, as well as the membership of the Quantum Internet Alliance for their broad support.
@@ Appendix A Derivation of Semiclassical Limit to Multiplexing
Here we derive rigorously the semiclassical upper-bounds to the multiplexing advantage introduced in section II. Throughout our paper we have defined the multiplexing advantage $m_{s}$, Mathematically, we can define $m_{s}$ through the expression
$R_{s|F}(M)=m_{s}R_{s|F}(1).$ (41)
If we are in the regime where almost all instantiations of the protocol fail, then each implementation may be considered a Poisson or geometric process. This enables us to interpret $m_{s}$ physically, in this limit, as the number of un-multiplexed implementation of the $s$-qubit quantum protocols that would need to run in parallel to achieve the same rate as a single instantiation of the $M$-fold multiplexed protocol.
We now begin with the derivation of the multiplexing advantage to the single-qubit rate to prepare or transmit a single qubit of fidelity $F$, which we provided a semiclassical limit for in Eq. 6. Rearranging Eq. 41 for $s=1$, we find
$m=\frac{R_{p_{M}|F}(M)}{R_{p|F}(1)}.$ (42)
Here $p$ is the probability that an un-multiplexed attempt to generate entanglement or remotely prepare a state succeeds, and $p_{M}$ is the probability that each attempt of the multiplexed attempts succeed. We can take $p_{M}=p$ to give an upper limit to $m$, which we assume henceforth. To reiterate, we also assume that the classical multiplexing does not negatively impact the maximum achievable fidelity for the protocol, as is the case e.g. for temporal multiplexing, up to a practical limit imposed by the physics e.g. how many pulses can fit in a beamline without overlap, how many spatial or spectral modes are accessible for multiplexing.
We now recast our un-multiplexed rate $R_{p|F}(1)$ in terms of expectation values $\tau_{p}$: the time to achieve the first (and here, only) success, which is a random variable. We assume this to be a Poisson process as is done in Ref. Davies et al. (2024). In this way for the unmultiplexed protocol, we can write
$R_{p|F}(1)=\frac{1}{\mathbb{E}(\tau_{p})(\tau_{1}+\tau_{c})}$ (43)
with $\tau_{1}$ the duration of a single attempt’s propogation time and $\tau_{c}$ the duration of any additional classical communication needed to complete the attempt.